OpenStudy (anonymous):

revolve the region bounded by x axis , y = sqrt ( x+ 4) , and x=0 about y axis using shell method. compare it to the disc method

OpenStudy (dumbcow):

\[=2 \pi \int\limits_{0}^{4}x \sqrt{x+4} dx\]

OpenStudy (anonymous):

no thats not the region

OpenStudy (anonymous):

the region is bounded by y = sqrt ( x + 4) , the y axis, its in quadrant 2

OpenStudy (anonymous):

so using that integral , its from -4 to 0

OpenStudy (dumbcow):

yes but we are revolving around y-axis so it covers both quadrants symmetrically and this way you get a positive volume

OpenStudy (anonymous):

it isnt symmertric though

OpenStudy (dumbcow):

x represents radius of 3-d figure, as you move away from y-axis the radius goes from 0 to 4

OpenStudy (anonymous):

the area from 0 to 4 is not equal to the area from -4 to 0

OpenStudy (anonymous):

how about this, revolve region bounded by y = sqrt x, x = 4, axis, about the line y = -4

OpenStudy (dumbcow):

oh i see, ok think of you reflected the function over y-axis replace x with (-x) f(x) = sqrt(4-x) i didn't do that in the intergral

OpenStudy (anonymous):

yes, that will do it,

OpenStudy (dumbcow):

sorry

OpenStudy (anonymous):

im trying to avoid these cases. are there other situations i can make up,

OpenStudy (anonymous):

no you did fine :)

OpenStudy (anonymous):

i dont seem to have a problem with disc method, since we have pi*r^2, but the shell method does produce negative

OpenStudy (dumbcow):

i guess always try to reflect onto positive quadrant when possible with volumes you want to integrate over positive values

OpenStudy (dumbcow):

make radius positive in shell method then adjust f(x) accordingly

OpenStudy (anonymous):

but i think the disc method in that case works

OpenStudy (anonymous):

using horizontal slices, we have x = y^2 - 4

OpenStudy (anonymous):

for y = sqrt ( x + 4

OpenStudy (dumbcow):

yeah discmethod \[=\pi \int\limits_{0}^{2}(y^{2}-4)^{2}dy\]

OpenStudy (anonymous):

so we get integral pi (

OpenStudy (anonymous):

pi integral ( 0 - (y^2-4))^2

OpenStudy (anonymous):

from -4 to 0

OpenStudy (dumbcow):

i prefer disc method

OpenStudy (anonymous):

oh my bad

OpenStudy (anonymous):

0 to 2